OpenStudy (anonymous):

1. If I know the lengths of two sides of a right triangle, how do I find the third? 2. Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? 3. Could I find the two missing angle measures if I know some of the side lengths of a right triangle? 4. What makes a triangle a “special” right triangle? How can special right triangles help me find side lengths?

OpenStudy (anonymous):

1. Yes, by using Pythagoras Theorem. \[a=\sqrt{b^2+c^2}\] 2. Yes, by using trigonometric functions.

OpenStudy (anonymous):

3. Again, you can find the angles by using trigonometric functions. \[\sin \theta= \frac{ \perp }{ hyp }\] \[\cos \theta= \frac{ base }{ hyp }\] \[\tan \theta= \frac{ \perp }{ base }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Special right triangles have these fixed angles: 1) 45-45-90 2) 30-60-90